Exercise 2.121 Let R be a ring. Definition 2.121.1 The center of R, written Z(R), is defined to be the set {re R | rx = xr fo #1 & 2
Answer #1

- 0 Z(R) is subring of R 0x = 0 = xo H NH? =) OE Z(R) Z(R) & ¢ let a, b EZER then (a - b)x= on-bn = xq- rb: x 19-b) - 9-b E Z